### v1

# RecipeGen: A Step-Aligned Multimodal Benchmark for Real-World Recipe Generation

Abstract.
Creating recipe images is a key challenge in food computing, with applications in culinary education and multimodal recipe assistants. However, existing datasets lack fine-grained alignment between recipe goals, step-wise instructions, and visual content. We present RecipeGen, the first large-scale, real-world benchmark for recipe-based Text-to-Image (T2I), Image-to-Video (I2V), and Text-to-Video (T2V) generation. RecipeGen contains 26,453 recipes, 196,724 images, and 4,491 videos, covering diverse ingredients, cooking procedures, styles, and dish types. We further propose domain-specific evaluation metrics to assess ingredient fidelity and interaction modeling, benchmark representative T2I, I2V, and T2V models, and provide insights for future recipe generation models. Project page is available at https://wenbin08.github.io/RecipeGen.
1. Introduction
“Food is the paramount necessity of the people.”
Generating illustrated or video-based instructions has attracted increasing attention as an effective way to guide users through complex, step-by-step tasks. Whether assembling furniture, repairing devices, or executing cooking procedures, visual representations—especially temporal ones—can significantly reduce user confusion and improve task success rates. One particularly compelling application is Recipe Generation, where step-by-step illustrations enhance the interpretability and accessibility of cooking instructions. Existing studies in recipe image generation can be categorized into creating dish images from the recipe title and ingredients (H. Lee et al., 2020), from recipe steps and ingredients (Pan et al., 2020; Liu et al., 2023; Han et al., 2020), or conditioned on a reference dish image (Chhikara et al., 2024; Dosovitskiy, 2020; Wang et al., 2020, 2022). While these approaches have shown promise, they often struggle to generate intermediate steps with clarity and consistency, which leads to confusion and misinterpretation. A primary reason for this limitation lies in existing datasets (Marın et al., 2021; Bień et al., 2020; Batra et al., 2020), which typically only include final images of the dish, lacking detailed snapshots that capture essential transitions between steps.
To advance research in recipe generation, we present a novel multimodal dataset named RecipeGen specifically designed for both recipe image and video generation tasks. Our dataset spans a broad spectrum of cuisines, cooking techniques, and dish categories, with meticulous step-level annotations comprising detailed textual descriptions paired with corresponding visual content. It consists of two well-aligned subsets: (1) a large-scale text-image collection featuring per-step visual illustrations, and (2) a curated text-image-video subset that includes step-aligned cooking videos alongside manually selected keyframes for each step. This dual-modality design enables unified training paradigms and facilitates cross-modal evaluation of generative models under both static and dynamic visual supervision.
Cooking procedures inherently involve complex ingredient transformations and intricate inter-ingredient interactions, posing significant challenges to maintaining semantic and visual consistency across steps. Furthermore, existing evaluation metrics largely fail to capture these fine-grained dynamics and cross-step coherency. To fill this gap, we propose a comprehensive evaluation framework consisting of three novel metrics: Cross-Step Consistency, which quantifies the visual coherence between consecutive steps; Ingredient Accuracy, which measures the fidelity of generated ingredients relative to ground-truth annotations; and Interaction Faithfulness, which assesses the correctness of depicted ingredient interactions in accordance with the described cooking actions. Together, these metrics offer a more task-specific and holistic assessment of recipe generation quality.
Our contributions are threefold:
High-quality step-aligned multimodal data. RecipeGen includes both text-image and text-image-video subsets, supporting a variety of generation tasks such as text-to-image, image-to-video, and text-to-video. It offers detailed, step-level textual and visual annotations with minimal non-instructional content, enabling fine-grained semantic and visual understanding of complex cooking procedures.
Large-scale and diverse dataset coverage. RecipeGen encompasses a wide array of ingredients, cooking techniques, and regional cuisines, supporting the development of robust, generalizable recipe understanding and generation models.
A novel and comprehensive evaluation suite. We introduce a hierarchical evaluation protocol with three new metrics—Cross-Step Consistency, Ingredient Accuracy, and Interaction Faithfulness—that better capture procedural correctness and semantic-visual alignment in generated recipes.
2. Related Work
Food is closely related to people’s lives, and with the development of human society, diets have diversified. To facilitate the management of human life and health, the field of food computing (Min et al., 2019b) has emerged. Academically, topics such as food segmentation (Lan et al., 2023; Yin et al., 2023), food recognition (Min et al., 2023; Jiang et al., 2020; Zhang et al., 2023), food recommendation (Min et al., 2019a; Wang et al., 2021), food reasoning (Zhou et al., 2024) and recipe image generation have gained significant attention.
Food Computing Dataset. There are several existing datasets for food computing tasks. For instance, Food-101 (Bossard et al., 2014), VireoFood-172 (Chen and Ngo, 2016), Recipes242k (Rokicki et al., 2018). However, these datasets are primarily used for fine-grained tasks such as food or ingredients classification, nutritional analysis, calorie calculation, etc. In addition to these information, FoodEarth (Zhou et al., 2024) utilizes LLMs to develop a comprehensive dataset for food and nutrition analysis. Similarly, Recipe1M (Marın et al., 2021) provides recipe instructions linked to each final dish image. However, neither dataset includes images for each individual recipe step. While the recipe section of VGSI (Zhou et al., 2018) offers step images for each instruction, many of these images are presented in a comic style. To address this gap, we propose the RecipeGen Benchmark, which is the first dataset that provides a real-world step-image recipe dataset.
3. RecipeGen Benchmark
| Dataset | #Recipes | Images | Videos | Modalities | Tasks | Source | Data Faithfulness | |
| Goal (GF) | Step (SF) | |||||||
| VGSI-Recipe (Yang et al., 2021) | 1,157 | 6,417 | 0 | T, I | T2I | wikiHow | 27.35 | 29.03 |
| YouCook2 (Zhou et al., 2017) | 2,000 | 0 | 2,000 | T, V | VU | YouTube | 27.77 | 29.93 |
| RecipeGen | 26,453 | 196,724 | 4,491 | T, I, V | T2I, I2V, T2V | Douguo, YouTube, YouCook2 | 28.94 | 30.31 |
The diversity of human culinary practices, developed over centuries, has led to a vast array of ingredient combinations and cooking methods, posing challenges for constructing comprehensive recipe step-image datasets. There are few existing datasets that cover a wide range of ingredients and interactions, and step counts are often limited.
To address the limitations of existing recipe datasets—such as limited step coverage, narrow ingredient diversity, and a predominant focus on Western cuisine—we introduce the RecipeGen Benchmark, a large-scale multimodal dataset containing both text-image and text-image-video recipes. RecipeGen comprises 26,453 curated recipes, paired with 196,724 step-aligned images and 4,491 cooking videos, covering a wide spectrum of regional dishes and culinary styles. The dataset is constructed from user-uploaded platforms, web-crawled resources, and existing datasets, offering a robust foundation for evaluating Text-to-Image (T2I), Image-to-Video (I2V), and Text-to-Video (T2V) generation models. Unlike previous datasets such as VGSI-Recipe (Yang et al., 2021) and YouCook2 (Zhou et al., 2017), which primarily focus on Western cuisine, RecipeGen captures a broader spectrum of culinary traditions. For instance, the same ingredient—beef—may be pan-fried as a steak in Western recipes, but slow-braised with root vegetables in East Asian dishes. By incorporating these regional variations, RecipeGen facilitates a greater understanding of diverse ingredient treatments and cross-cultural cooking strategies.
3.1. Construction Procedure for Image-Text Recipes
To ensure RecipeGen’s diversity, we curate 158 keywords across different cuisines (e.g., Cantonese, Mexican), dish types (e.g., main courses, desserts), and cooking techniques (e.g., stir-frying, roasting), gathering 21,944 recipes.
Keyword Selection and Coverage. A total of 158 keywords covering 158 different cuisines, cooking methods, and dish types are defined to ensure broad and structured coverage. Fig. 2 illustrates their distribution. Based on these keywords, we collect a total of 29,026 recipes.
Quality Control. To maintain consistency and clarity, we implement a quality control process to address common issues like incomplete instructions or overly detailed steps. This process involved: (1) Removing recipes with low quality (including lack images or steps, mismatched numbers of images and steps, vague directions 111For example, some steps in the recipe only contain phrases like “see as images” without providing clear or actionable instructions.). In this step, we delete 4,978 recipes; (2) Using GPT-4o (Achiam et al., 2023) to clean and refine recipes by eliminating irrelevant content 222For instance, some users include exclamatory phrases such as “This tastes amazing!” or suggestions like “Using ingredient A instead of B works better.”, merging redundant steps, and summarizing for coherence. And then, we remove recipes where the LLM’s hallucination issues caused a single step to be split into two, as well as recipes with incorrect output formats. A total of 2,104 recipes were deleted. After Quality Control, the number of recipes is 21,944 and the number of images is 139,872; (3) Computing data faithfulness and using human check to insure the quality. The result of metrics is in Tab. 1.
Metrics Computation and Human Check. To assess dataset quality, we use three metrics. Goal Faithfulness (GF) measures the CLIP score between the final image and its caption, reflecting dish completeness. Step Faithfulness (SF) computes the CLIP score between each image and its corresponding step. As shown in Tab. 1, our RecipeGen dataset outperforms VGSI-Recipe and YouCook2 across all metrics, with GF of 28.94, SF of 30.31. In addition to evaluating the dataset quality using step faithfulness and goal faithfulness metrics, we also conduct a human review. Specifically, annotators are asked to verify three aspects: (1) whether the recipe matches its dish name, (2) whether each step corresponds correctly to its associated image, and (3) whether the images within the same recipe exhibit visual continuity. We randomly sample 2,000 recipes, covering more than 12,000 images for this review. Among these samples, we identify 12 instances where the dish name was mislabeled (e.g., “labeling dumplings” as “steamed buns”), and 7 instances where all the textual steps were generic and lacked specific ingredient descriptions (e.g., using phrases like “as shown in the image” or “ready to serve”). As well as 6 recipes in which no pictures existed and 17 cases that lacked pictures corresponding to some of the steps.
3.2. Construction Procedure for Image-Text-Video Recipes
For the video data, the sources primarily consist of web-crawled videos and a curated subset of the YouCook2 (Zhou et al., 2017) dataset. In total, RecipeGen includes 4,491 recipes paired with instructional videos.
Problem of Existing Dataset. YouCook2 is a dataset collected from YouTube that contains 2,000 cooking videos with temporal annotations for each step 333We only utilize the training and validation sets, as these are the only subsets that provide detailed step-wise annotations necessary for our processing. In total, these two splits contain 1,774 recipe videos., including start and end timestamps, along with corresponding video segments. However, the dataset presents several limitations for step-level recipe generation and food-centric modeling. First, as shown in Fig. 1, because YouCook2 was designed for general video captioning rather than recipe-specific tasks, and its videos are often sourced from lifestyle vloggers, the resulting content is typically human-dominant rather than food-dominant, with limited focus on ingredients and cooking actions. To quantify this issue, we use YOLOv8 (Jocher et al., 2023) to estimate the temporal proportion of frames containing people. Sampling one frame every 10 frames and applying a confidence threshold of 0.5 for the “person” label, we find that the average proportion of person-visible duration across the dataset is 46.58%, with 850 videos having more than 50% of their duration focused on people. This significantly reduces the visual relevance for food-centric tasks. Second, because each step is represented as a video segment rather than an individual image, YouCook2 lacks explicit step-level image modality, limiting its applicability for tasks such as step-wise image generation or vision-language alignment at the step level.
Quality Control for YouCook2. To adapt YouCook2 for recipe video generation, we applied several preprocessing steps. First, we used YOLOv8 (confidence ¿ 0.5) to detect humans in sampled frames (1 every 10 frames). Videos with over 50% person presence were discarded; for the rest, person-containing frames were removed and clips reassembled. This yielded 924 videos, reducing person-visible content to just 0.98%, ensuring focus on cooking actions and ingredients.
Second, since each annotated step in YouCook2 corresponds to a continuous video segment rather than a single representative image, we extract a keyframe for each step to enable image-based modeling. Specifically, for each step, we compute the CLIP similarity between the step’s textual description and all frames within the corresponding video segment. The frame with the highest CLIP similarity score is selected as the keyframe representing that step. Finally, a total of 924 recipes are retained.
Quality Control for Web-Crawled Data. We expanded the dataset by collecting recipe videos from Douguo444www.douguo.com and YouTube555www.youtube.com. Douguo offers aligned step-by-step instructions, images, and videos, requiring no additional processing. YouTube lacks explicit step annotations but provides auto-generated “key moments,” which we extract via lazy loading during crawling. As the associated thumbnails are often low-resolution, we enhance quality by sampling frames around each timestamp and selecting the one with the highest SSIM (Wang et al., 2004) score to replace the thumbnail, ensuring visual clarity and semantic consistency.
To further improve data quality, we apply a two-stage filtering process:
1) Automated Text Cleaning: A GPT-based method removes non-informative phrases and irrelevant labels common in user-generated content.
2) Manual Alignment: We manually correct mismatches between cleaned text and images to maintain accurate step-image pairs. After this process, a total of 3,567 recipes are retained.
3.3. Metrics
Cooking is an inherently complex, multi-step process that involves sequential actions, ingredient interactions, and temporal dependencies. To faithfully evaluate model performance in such a setting, we design metrics along two primary dimensions, , as summarized in Tab.2:
Step-Level Faithfulness. At the step level, each generated image is evaluated for its alignment with the corresponding textual instruction in two aspects: ingredient accuracy, which measures whether all required ingredients are correctly and completely depicted, and interaction accuracy, which assesses whether the visualized ingredient interactions match the semantics described in the step. We categorize interaction types into five visual patterns: (1) Mix up—multiple ingredients appear together while retaining their original form (e.g., tomato and scrambled egg); (2) Blend—ingredients are fully fused into a new form with no distinct identity (e.g., red bean paste mixed into dough); (3) Put on—one ingredient is simply placed atop another without merging (e.g., cheese on pizza); (4) Cover—one ingredient fully wraps or encloses another (e.g., dumpling wrapper enclosing filling); and (5) No relationship—no interaction occurs between ingredients, typically during early-stage preparations such as cutting or washing. Detailed example is in Fig.1.
Cross-Step Consistency. Beyond individual steps, we assess Cross-Step Consistency by measuring whether visual attributes of the same ingredients are maintained across steps and whether the total number of generated steps aligns with the ground truth, using visual distance and step count difference as indicators.
| Metric | Description |
Goal
Faithfulness |
CLIP (Radford et al., 2021) similarity between the final image and last-step caption, measuring alignment with the overall goal. |
Step
Faithfulness |
Assesses each image’s alignment with its step caption using CLIP. |
Ingredient
Accuracy |
Uses GPT-4o (Achiam et al., 2023) to extracts step-wise ingredients and finetune Qwen2.5-VL (Bai et al., 2025) for three epochs to predict them from generated images. The trained model serves as a VQA model to detect the presence of expected ingredients, with the ratio of correctly detected to ground-truth ingredients as accuracy. |
Interaction
Faithfulness |
Based on different interaction types—Mix up, Blend, Put on, Cover, and No relationship—we design a standardized prompt that guides GPT-4o to generate one-sentence captions grounded in visible features such as color, texture, spatial layout, and perceived fusion. We then compute the CLIP similarity score between each generated image and its corresponding caption to evaluate interaction consistency. |
Cross-Step
Consistency |
Based on StackedDiffusion (Menon et al., 2024) and DINOv2 (Caron et al., 2021), uses distance and step count difference to assess visual and numerical consistency. |
3.4. Dataset Statistics
Overall, our proposed RecipeGen Benchmark has four notable features:
Broad Step Distribution and Ingredient Diversity: The dataset includes recipes with 2 to 25 steps, with an average of 7.4 steps, supporting the modeling of long-range semantic relationships. RecipeGen provides an average of 9 ingredients per recipe (including seasonings), capturing a rich variety of ingredients and interactions.
Variety of Cooking Styles: The dataset encompasses a wide range of cooking styles enhanced by numerous unique keywords, making it versatile across different culinary processes.
High-quality step-aligned multimodal content: Our dataset contains cooking videos with a high proportion of effective instructional frames, significantly reducing non-informative host appearances. Each video is accompanied by detailed, step-by-step textual instructions and corresponding diagrammatic illustrations, which together enable a fine-grained understanding of the cooking process.
Comprehensive evaluation metrics: We design evaluation metrics that capture both the overall visual quality and fine-grained semantic accuracy. These metrics assess not only the coherence and realism of the generated content at a high level, but also verify the presence, correctness, and interactions of individual ingredients throughout each recipe step.
Real-World Representativeness: Collected from various users, RecipeGen consists entirely of real-world recipes and closely resembles the types of instructions that users might upload in practical scenarios. Quality control using GPT-4o ensures that steps are trustworthy, reflecting actual cooking practices.
4. Conclusion
In this work, we present RecipeGen, the first step-aligned, multimodal benchmark for real-world recipe generation. By providing richly annotated text, image, and video data, RecipeGen addresses the lack of comprehensive datasets and metrics in food computing. We believe that our benchmark can foster further advancements in food computing, particularly in real-world, interactive culinary applications.
5. Usage and License
All data used in the RecipeGen benchmark is collected from publicly available sources. RecipeGen is released under the Creative Commons Attribution-NonCommercial 4.0 International (CC BY-NC 4.0) license. This license permits sharing and adaptation of the dataset for non-commercial purposes, provided appropriate credit is given. Commercial use is strictly prohibited under the terms of this license. The dataset is publicly available under the CC BY-NC 4.0 license at HuggingFace.